A hotel claims that ​85% of its customers are very satisfied with its service. Complete parts a through d below based on a random sample of 7 customers.
What is the probability that less than 5 customers are very​ satisfied?

Found 2 solutions by ikleyn, Boreal:
Answer by ikleyn(52878)

.

No parts a through d in the post.


You missed them somewhere.



Answer by Boreal(15235)

Fewer than 5 customers.
Do probabilities of 5,6,7 and subtract those 3 probabilities from 1.
for 5: 7C5*0.85^5*0.15^2=0.2097
for 6: 7*0.85^6*0.15=0.3960
for 7: 0.85^7=0.3206
they add to 0.9263
so the answer is the complement or 0.0737